# Prospects for High-Frequency Gravitational-Wave Detection with GEO600

### v2

@@ Abstract
Current ground-based interferometers are optimized for sensitivity from a few tens of Hz to about 1 kHz. While they are not currently utilized for GW detection, interferometric detectors also feature narrow bands of strong sensitivity at higher frequencies where the sideband fields created by a GW are resonantly amplified in the optical system. For certain interferometer configurations, small changes to system parameters allow the narrow band of high sensitivity to be scanned over a much larger range of frequencies, potentially enabling broadband detection at high frequencies. In this paper, we investigate whether simply modifying the detuning angle of the signal-recycling mirror of the GEO600 interferometer can make this experiment sensitive to GWs in the kilohertz frequency range. We compute the strain sensitivity for GEO600 across a frequency range from several kHz to tens of kHz for various detuning angles. We also show that LIGO cannot attain the same effect assuming that the optical components are not changed due to the narrow band response of the Fabry-Perot cavities. We then calculate the sensitivity of GEO600 to various proposed high-frequency GW sources and compare it to the sensitivity of other ground-based detectors.
@@ I Introduction
Detecting GWs at frequencies in the tens of kHz holds the potential to unveil signatures of new physics such as very light compact object binaries or ultralight boson clouds [42, 43]. Sub-solar mass (SSM) compact objects could be astrophysical objects like BHs created through an unconventional formation channel [44, 45, 46], a primordial BH [47, 48, 49], or some other exotic compact object [50, 51, 52, 53]. If multiple SSM objects form a binary, it will be detectable by searching for its effects with a high-frequency GW detector [54, 55]. Additionally, isolated spinning black holes can form boson clouds via the superradiance effect if there is some unknown ultralight boson cloud [56, 57]. Detection of the GWs emitted by the boson clouds could be evidence of new physics. The quantum chromodynamics (QCD) axion [58, 59, 60, 61] and the dark photon [62, 63, 64] are two prominent ultralight bosons that could have astrophysical effects in GWs [65, 66, 67]. The LVK collaboration has actively searched for GWs from SSM binaries [68, 69, 70, 71, 72] and superradiant boson clouds in their data [73, 74, 75]. SSM binaries and superradiant boson clouds may generate GWs with characteristic frequencies above the sensitivity bands of current generation detectors [76]. In addition to superradiance, ultralight bosons could be bound to the gravitational potential of a NS [77, 78] or BH which would allow indirect detection via fifth-force in GW signals [79, 80, 81] or in binary pulsar observations [82, 83, 84, 85]. Outside of GW observations, ultralight bosons and SSM compact objects have been searched for with both direct and indirect experiments; however, no evidence has been found [86, 87, 88, 89, 90, 91, 92, 93]. Lastly, enhancing the high-frequency sensitivity of gravitational-wave detectors expands the range of astrophysical phenomena that can be probed through data analysis [94].
@@
The rest of this paper is organized as follows. In Sec. II, we discuss the underlying detector physics that motivated the investigation of GEO600 as a high-frequency GW detector. We describe our methods for computing strain sensitivities for detuned configurations of GEO600. We discuss the resulting sensitivities of different interferometer configurations and compare them to other ground-based interferometric detectors. In Sec. III, we introduce the high-frequency astrophysical sources and describe their GW waveforms and properties. We discuss how the SNR is calculated and the resulting ability of each interferometer to detect the sources. In Sec. IV, we conclude our work and discuss future research. The data used in this paper can be downloaded from GitHub [100].
@@ II.1 Underlying Physics
To improve the sensitivity of a GW detector, one can either weaken the background noise sources or amplify the magnitude of the detector’s response to an incident GW. GEO600’s potential as a high-frequency GW detector originates in the latter option, namely through the resonant enhancement of GW-induced sideband fields within the interferometer [101]. Resonant amplification of the sideband fields in turn amplifies the signal power at the detection port, boosting the overall response. GEO600 features two mirrors which reflect laser light departing the Michelson arms back into the MI, thus forming effective cavities with the north and east Michelson arm mirrors (shown in Fig. 1). The power-recycling cavity (PRC), created by the power-recycling mirror (MPR) and MI, increases the sensitivity of the interferometer by building up the cavity light’s intensity and distance traveled [102]. The signal-recycling cavity (SRC), created by the MSR and MI, allows the signal sidebands to resonate at specific Fourier frequencies, denoted resonant frequencies [103, 104, 105].
@@
For this work, we primarily target the frequency which appears for n=0n=0111The resonant frequencies for $|n|>0$ are less advantageous for detection because the GW-induced phase delay in the MI is diminished for frequencies greater than $f_{\mathrm{fsr}}=c/2L_{\mathrm{src}}$ (see Appendix A).. The detuning angle $\phi$ is related to the MSR displacement $\delta l$ by $\phi=k\delta l$, where $k=2\pi\nu/c$. Therefore, by finely tuning the MSR location, we can scan the SRC resonant frequencies, and thus the resonant peaks in the GEO600 detector response, over a broad frequency range [106]. Detuning the MSR allows for lowest-order resonant peaks in the detector response function to occur as high as 62.5 kHz.
@@
Unlike GEO600, the LIGO detector response displays two distinct resonant features, but we will see that they are not well suited for broadband high-frequency detection with current mirror properties. The first occurs at multiples of the FSR of the Fabry-Perot arms (37.5 kHz), but it is unaffected by the detuning of the MSR, and thus cannot be scanned with this procedure. One might envision detuning the Fabry-Perot mirrors themselves, but this would drastically reduce the DC power circulating in the arms (that is, intentionally moving the cavities away from their locked position), heavily attenuating the available power for sideband generation.
@@
The other feature can be scanned by detuning the SRC; however, it cannot reach into our target frequency band. From a optical resonance perspective, the Fabry-Perot arm cavities increase the effective SRC length by a factor of $\sim 2\mathcal{F}$, where $\mathcal{F}$ is the cavity finesse. This limits the maximum first-order resonance frequency to $\sim 200$ Hz with current LIGO mirror transmissivities; all higher-order resonant frequencies of the SRC are washed out by the high-finesse Fabry-Perot arms.
@@
As it turns out, though, this feature is not produced exclusively by optical resonance behavior. In fact, in the small detuning angle regime, the frequency of optimal sensitivity (in the $\sim 100$ Hz region) scales as $\sqrt{\phi P_{\mathrm{laser}}}$, indicative of the presence of an optical spring effect [107]. The optical spring creates sharp features in the aLIGO sensitivity function for small detuning angles. However, as the detuning angle increases, the optical spring frequency increases, but its linewidth increases at a faster rate, resulting in a broader feature. The optical spring frequency is also limited to around 200 Hz with the current properties of LIGO optics, so neither the optical resonance nor optical spring effects can generate improvements to the LIGO sensitivity function outside its nominal frequency band simply by detuning the signal recycling mirror. Thus, a GEO600-like detector geometry — namely, one featuring shorter arms (and thus a higher FSR) and a lack of Fabry-Perot cavities — is more favorable for broadband high-frequency detection via the procedure we have described.
@@ II.2 Methods
Beam splitter
$0.513872$ $130$ ppm $9.3$ kg
Central East mirror
$13$ ppm $130$ ppm $5.6$ kg
Central North mirror
$13$ ppm $130$ ppm $5.6$ kg
Far East mirror
$8.3$ ppm $130$ ppm $5.6$ kg
Far North mirror
$8.3$ ppm $130$ ppm $5.6$ kg
Power-recycling mirror
$900$ ppm $130$ ppm $2.92$ kg
Signal-recycling mirror
$0.09995$ $50$ ppm $2.92$ kg
@@
To compute $S_{h}(f)$, we employ the Finesse 3.0 package [112, 113], which we use to extract the frequency-domain transfer functions $\partial P/\partial n_{i}$ [^2] Finesse features a method to compute the quantum noise at any optical port, so for quantum noise in particular, we directly extract the noise at the detection port. and $\partial P/\partial h$ from a model of GEO600. Notably, the transfer functions from a GW to an observable are computed in Finesse under the assumption that the GW is incident from directly above the interferometer. Denoting the transfer function to the dark port power calculated by Finesse as $\frac{\partial P}{\partial h}\bigr|_{\texttt{Finesse}}$, the sky-averaged detector response function is therefore equal to
@@
Figure 2: Strain sensitivity $A_{h}(f)$ of GEO600 at different detuning angles of the MSR, as well as a scanned sensitivity curve (black), which is a summation of all possible MSR detuning angles. As the detuning angle of the MSR is increased from $\phi=0$ (tuned configuration for GEO600), the detection region becomes increasingly narrow-band, in that the frequency bandwidth stays constant while the SRC resonant frequency increases. The scanned sensitivity curve is plotted to show the possible values that can be achieved across the kHz frequency range by shifting the MSR.
Figure 3: Strain sensitivity $A_{h}(f)$ of the aLIGO design at different detuning angles of the MSR. When the angle of the MSR is decreased from $\phi=90$ (tuned configuration for aLIGO that corresponds to ‘signal extraction’ [115, 106]), a region of strong sensitivity gradually shifts to lower frequencies in the neighborhood of $f\sim\mathcal{O}(100~\text{Hz})$, while its bandwidth simultaneously narrows. Note that we use $\phi=1$ to represent the ‘signal recycling’ limit, as setting $\phi=0$ exactly corresponds to a configuration completely dominated by classical noise. There is no enhancement in sensitivity at frequencies in the kHz from aLIGO detuning the MSR. The peaks in sensitivity occurring at integer multiples of $37.5$ kHz (FSR) are created by the Fabry-Perot cavities and are independent of the MSR detuning angle. Note that in this plot we are using current LIGO mirrors — it is possible that changing the MSR or TM transmissivities then LIGO could have improved high-frequency sensitivity.
Figure 4: Strain sensitivity $A_{h}(f)$ of GEO600 tuned (purple), anti-tuned (gray), and scanned (dashed), which is a summation of all possible MSR detuning angles. The aLIGO design sensitivity (red) [116] and Cosmic Explorer’s design sensitivity (light blue) [116] are also shown. GEO600’s scanned sensitivity curve shows the possible frequencies in the kHz range where it would have an advantageous narrow band detection to that of Cosmic Explorer and aLIGO.
Note that when we plot detector strain sensitivities in this work (e.g., Figs. 2, 3, 4, and 6), we assume a fiducial GW sky location of $\theta_{\hat{n}}=15$ and $\phi_{\hat{n}}=0$, as is typically done in the CE community [117, 118].
@@ II.3 Sensitivities
To assess the potential advantages of using GEO600 as a detection instrument for each source, we compared the signal-to-noise ratio (SNR) of the aLIGO design, GEO600, and GEO600 detuned to various frequencies. The sky-averaged strain sensitivities of these interferometer configurations, as well as that of Cosmic Explorer’s [16] for further comparison, are shown in Fig. 4. GEO600’s scanned sensitivity curve shows the possible frequencies in the kHz range where it would have an advantageous narrow band detection to that of Cosmic Explorer and aLIGO.
@@ III.1.1 GW Strain Model
Figure 5: Left: contour lines corresponding to a SNR of eight for GWs sourced from both vector (solid) and scalar (dashed) boson clouds. The contour lines correspond to the strain sensitivities of aLIGO (blue), tuned GEO600 (orange), and GEO600 with different MSR detuning angles. The shaded gray region corresponds to the parameter space where the superradiance condition fails. An SNR of eight can be achieved at frequencies as large as $31$ kHz, with the highest frequency obtained at a detuning angle of $\phi=45\degree$ (not plotted). Right: SNR of GWs sourced from both vector (solid) and scalar (dashed) boson clouds across a range of boson rest energies for a 0.3 $M_{\odot}$ BH. The SNR curves are calculated with the same strain sensitivities as the left plot and have the same corresponding colors. The dotted black line represents where the SNR is eight. The sharp increase in SNR in aLIGO at 37.5 kHz (FSR) arises due to sideband resonance in its Fabry-Perot cavities, while the kink in the SNR function on the right plot around 14 kHz corresponds to the sudden drop-off in integration time when $\tau_{\text{GW}}$ falls below one week. All GWs are sourced from within the galaxy, being simulated from a distance of $30$ kPc.
@@
where $\omega_{\mu}=\mu/\hbar$, $m=1$, and $\Omega_{\mathrm{BH}}(0.7)=1.38\times 10^{5}\!\left(\!\frac{0.3M_{\odot}}{M}\!\right)\text{s}^{-1}$. Rearranging this superradiance condition constrains the rest energy of bosons surrounding a BH of mass $M$ to be $\mu<9.10\times 10^{-11}\!\left(\!\frac{0.3M_{\odot}}{M}\!\right)\text{eV}$. For boson clouds with boson rest energies in the target range of $10^{-12}$ to $10^{-10}$ eV, the GW emission frequency is given by [124] as
$f_{\text{GW}}\approx 48750\,\text{Hz}\times\frac{\mu}{10^{-10}\text{ eV}}.$ (6)
@@
$\hskip-7.59995pth_{0}^{v}\approx 1.28\times 10^{-22}\!\left(\!\frac{M}{0.3M_{\odot}}\!\right)\!\!\left(\frac{\alpha}{0.2}\right)^{\!5}\!\!\left(\!\frac{0.03\,\text{Mpc}}{r}\!\right)\!\!\left(\!\frac{0.7-\chi_{f}}{0.1}\!\right)\!,$ (7)

### v3

@@ Abstract
Current ground-based interferometers are optimized for sensitivity from a few tens of Hz to about 1 kHz. While they are not currently utilized for GW detection, interferometric detectors also feature narrow bands of strong sensitivity at higher frequencies where the sideband fields created by a GW are resonantly amplified in the optical system. Small changes to system parameters allow the narrow band of high sensitivity to be scanned over a much larger range of frequencies. In this paper, we investigate whether simply modifying the detuning angle of the signal-recycling mirror of the GEO600 interferometer can make this experiment sensitive to GWs in the kilohertz frequency range. We compute the strain sensitivity for GEO600 across a frequency range from several kHz to tens of kHz for various detuning angles. We also show that LIGO cannot attain the same effect assuming that the optical components are not changed due to the narrow band response of the Fabry-Perot cavities. We then calculate the sensitivity of GEO600 to various proposed high-frequency GW sources and compare it to the sensitivity of other ground-based detectors.
@@ I Introduction
Detecting GWs at frequencies in the tens of kHz holds the potential to unveil signatures of new physics such as very light compact object binaries or ultralight boson clouds [128, 19]. Sub-solar mass (SSM) compact objects could be astrophysical objects like BHs created through an unconventional formation channel [91, 106, 126], a primordial BH [141, 76, 46], or some other exotic compact object [111, 100, 107, 93]. If multiple SSM objects form a binary, it will be detectable by searching for its effects with a high-frequency GW detector [127, 103]. Additionally, isolated spinning black holes can form boson clouds via the superradiance effect if there is some unknown ultralight boson cloud [29, 43]. Detection of the GWs emitted by the boson clouds could be evidence of new physics. The quantum chromodynamics (QCD) axion [114, 113, 137, 135] and the dark photon [67, 115, 60] are two prominent ultralight bosons that could have astrophysical effects in GWs [85, 27, 38]. The LVK collaboration has actively searched for GWs from SSM binaries [5, 6, 8, 109, 102] and superradiant boson clouds in their data [7, 131, 73]. SSM binaries and superradiant boson clouds may generate GWs with characteristic frequencies above the sensitivity bands of current generation detectors [61]. In addition to superradiance, ultralight bosons could be bound to the gravitational potential of a NS [65, 92] or BH which would allow indirect detection via fifth-force in GW signals [83, 84, 24] or in binary pulsar observations [130, 136, 125, 124]. Outside of GW observations, ultralight bosons and SSM compact objects have been searched for with both direct and indirect experiments; however, no evidence has been found [31, 22, 28, 10, 32, 49, 132, 96]. Lastly, enhancing the high-frequency sensitivity of GW detectors expands the range of astrophysical phenomena that can be probed [50].
@@
The rest of this paper is organized as follows. In Sec. II, we discuss the underlying detector physics that motivated the investigation of GEO600 as a high-frequency GW detector. We describe our methods for computing strain sensitivities for detuned configurations of GEO600. We discuss the resulting sensitivities of different interferometer configurations and compare them to other ground-based interferometric detectors. In Sec. III, we introduce the high-frequency astrophysical sources and describe their GW waveforms and properties. We discuss how the signal-to-noise ratio (SNR) is calculated and the resulting ability of each interferometer to detect the sources. In Sec. IV, we conclude our work and discuss future research. The data used in this paper can be downloaded from GitHub [88].
@@ II.1 Underlying Physics
To improve the sensitivity of a GW detector, one can either weaken the background noise sources or amplify the magnitude of the detector’s response to an incident GW relative to the response to perturbations in other channels. GEO600’s potential as a high-frequency GW detector originates in the latter option, namely through the resonant enhancement of GW-induced sideband fields within the interferometer [71]. Resonant amplification of the sideband fields in turn amplifies the signal power at the detection port, boosting the overall response. GEO600 features two mirrors which reflect laser light departing the Michelson arms back into the MI, thus forming effective cavities with the north and east Michelson arm mirrors (shown in Fig. 1). The power-recycling cavity (PRC), created by the power-recycling mirror (MPR) and MI, increases the sensitivity of the interferometer by building up the cavity light’s intensity and distance traveled [119]. The signal-recycling cavity (SRC), created by the signal-recycling mirror (MSR) and MI, allows the signal sidebands to resonate at specific Fourier frequencies, denoted resonant frequencies [101, 104, 78].
@@
For this work, we primarily target the frequency which appears for n=0n=0111The resonant frequencies for $|n|>0$ are less advantageous for detection because the GW-induced phase delay in the MI is diminished for frequencies greater than $f_{\mathrm{fsr}}=c/2L_{\mathrm{src}}$ (see Appendix A).. The detuning angle $\phi$ is related to the MSR displacement $\delta l$ by $\phi=k\delta l$, where $k=2\pi\nu/c$. Therefore, by finely detuning the MSR location, we can scan the SRC resonant frequencies, and thus the resonant peaks in the GEO600 detector response, over a broad frequency range [79]. Detuning the MSR allows for lowest-order resonant peaks in the detector response function to occur as high as 62.5 kHz.
@@
Unlike GEO600, the LIGO detector response displays both Fabry-Perot and signal-recycling resonant features, but we will see that it is not well suited for high-frequency detection with current mirror properties. The Fabry-Perot resonant feature occurs at multiples of the FSR of the Fabry-Perot arms (37.5 kHz), but it is unaffected by the detuning of the MSR, and thus cannot be scanned with this procedure. One might envision detuning the Fabry-Perot mirrors themselves, but this would drastically reduce the DC power circulating in the arms (that is, intentionally moving the cavities away from their locked position), heavily attenuating the available power for sideband generation.
@@
The signal-recycling resonant feature can be scanned by detuning the SRC; however, it cannot reach into our target frequency band. From a optical resonance perspective, the Fabry-Perot arm cavities increase the effective SRC length by a factor of $\sim 2\mathcal{F}$, where $\mathcal{F}$ is the cavity finesse. This limits the maximum first-order resonance frequency to $\sim 200$ Hz with current LIGO mirror transmissivities; all higher-order resonant frequencies of the SRC are washed out by the high-finesse Fabry-Perot arms.
@@
As it turns out, though, this feature is not produced exclusively by optical resonance behavior. In fact, in the small detuning angle regime, the frequency of optimal sensitivity (in the $\sim 100$ Hz region) scales as $\sqrt{\phi P_{\mathrm{laser}}}$, indicative of the presence of an optical spring effect [45]. The optical spring creates sharp features in the aLIGO sensitivity function for small detuning angles. However, as the detuning angle increases, the optical spring frequency increases, but its linewidth increases at a faster rate, resulting in a broader feature. The optical spring frequency is also limited to around 200 Hz with the current properties of LIGO optics, so neither the optical resonance nor optical spring effects can generate improvements to the LIGO sensitivity function outside its nominal frequency band simply by detuning the signal recycling mirror. Thus, a GEO600-like detector geometry — namely, one featuring shorter arms (and thus a higher FSR) and a lack of Fabry-Perot cavities — is more favorable for high-frequency detection via the procedure we have described.
For clarity, we define the interferometer terminology used throughout this paper. The tuned interferometer configuration corresponds to a MSR detuning angle of $\phi=0$ for GEO600 and $\phi=90$ for aLIGO. The anti-tuned configuration corresponds to a MSR detuning angle of $\phi=90$ for GEO600 and $\phi=0$ for aLIGO [^2] In practice for this configuration, we use $\phi=1$ since taking the limit as $\phi\rightarrow 0$ exactly corresponds to a configuration completely dominated by classical noise.. A detuned configuration corresponds to any MSR detuning angle between $0<\phi<90$ for GEO600 aLIGO. For GEO600, we use a scanned configuration to show the pointwise optimization in frequency space over MSR detuning angles. This scanned configuration is used solely as a visualization tool, as it is not a physically utilizable sensitivity itself.
@@ II.2 Methods
Beam splitter (BS) $0.513872$ $130$ ppm $9.3$ kg
Central East mirror (MCE) $13$ ppm $130$ ppm $5.6$ kg
Central North mirror (MCN) $13$ ppm $130$ ppm $5.6$ kg
Far East mirror (MFE) $8.3$ ppm $130$ ppm $5.6$ kg
Far North mirror (MFN) $8.3$ ppm $130$ ppm $5.6$ kg
Power-recycling mirror (MPR) $900$ ppm $130$ ppm $2.92$ kg
Signal-recycling mirror (MSR) $0.09995$ $50$ ppm $2.92$ kg
@@
To compute $S_{h}(f)$, we employ the Finesse 3.0 package [62, 41], which we use to extract the frequency-domain transfer functions $\partial P/\partial n_{i}$ [^3] Finesse features a method to compute the quantum noise at any optical port, so for quantum noise in particular, we directly extract the noise at the detection port. and $\partial P/\partial h$ from a model of GEO600. The numerical methods of Finesse have been validated and do not need low-frequency approximations [41]. The transfer functions from a GW to an observable are computed in Finesse under the assumption that the GW is incident from directly above the interferometer. Denoting the transfer function to the dark port power calculated by Finesse as $\frac{\partial P}{\partial h}\bigr|_{\texttt{Finesse}}$, the sky-averaged detector response function is therefore equal to
@@
Note that while we use sky averaging for SNR/detectability calculations, when we plot detector strain sensitivities in this work (e.g., Figs. 2, 3, 4, and 6), we assume a fiducial GW sky location of $\theta_{\hat{n}}=15$ and $\phi_{\hat{n}}=0$, as is typically done in the CE community [129, 59].
@@ II.3 Sensitivities
Figure 2: Strain sensitivity $A_{h}(f)$ of GEO600 at different detuning angles of the MSR, as well as a scanned sensitivity curve (black), which is the pointwise optimization over MSR detuning angles. As the detuning angle of the MSR is increased from $\phi=0$ (tuned configuration for GEO600), the detection region becomes increasingly narrow-band, in that the frequency bandwidth stays constant while the SRC resonant frequency increases. The scanned sensitivity curve is plotted to show the possible values that can be achieved across the kHz frequency range by shifting the MSR. Note that all strain sensitivities assume a GW sky location of $\theta_{\hat{n}}=15$ and $\phi_{\hat{n}}=0$.
Figure 3: Strain sensitivity $A_{h}(f)$ of the aLIGO design at different detuning angles of the MSR. When the angle of the MSR is decreased from $\phi=90$ (tuned configuration for aLIGO that corresponds to ‘signal extraction’ [44, 79]), a region of strong sensitivity gradually shifts to lower frequencies in the neighborhood of $f\sim\mathcal{O}(100~\text{Hz})$, while its bandwidth simultaneously narrows. Note that we use $\phi=1$ to represent the ‘signal recycling’ limit, as setting $\phi=0$ exactly corresponds to a configuration completely dominated by classical noise. There is no enhancement in sensitivity at frequencies in the kHz from aLIGO detuning the MSR. The peaks in sensitivity occurring at integer multiples of $37.5$ kHz (FSR) are created by the Fabry-Perot cavities and are independent of the MSR detuning angle. Note that in this plot we are using current LIGO mirrors — it is possible that changing the MSR or TM transmissivities then LIGO could have improved high-frequency sensitivity. Note that all strain sensitivities assume a GW sky location of $\theta_{\hat{n}}=15$ and $\phi_{\hat{n}}=0$.
Figure 4: Strain sensitivity $A_{h}(f)$ of GEO600 tuned (purple), anti-tuned (gray), and scanned (dashed), which is a summation of all possible MSR detuning angles. The aLIGO design sensitivity (red) [121] and Cosmic Explorer’s design sensitivity (light blue) [121] are also shown. GEO600’s scanned sensitivity curve shows the pointwise optimization in frequency space over MSR detuning angles. Note that all strain sensitivities assume a GW sky location of $\theta_{\hat{n}}=15$ and $\phi_{\hat{n}}=0$.
@@
To assess the potential advantages of using GEO600 as a detection instrument for each source, we compared the SNR of the aLIGO design, GEO600, and GEO600 detuned to various frequencies. We include sky averaging in SNR calculations to account for the reduced sensitivity at certain GW incidence angles where the detector response approaches zero, as shown in FIG. 7. In App. A, we discuss the sky-averaging in SNR calculations for the high-frequency antenna response of GEO in more detail (c.f. Fig. 7). The strain sensitivities of these interferometer configurations, as well as that of Cosmic Explorer’s [120] for further comparison, are shown in Fig. 4. GEO600’s scanned sensitivity curve shows the ASD optimized pointwise over MSR detuning angles at every possible frequency in the kHz range.
@@ III.1.1 GW Strain Model
Figure 5: Left: contour lines corresponding to a SNR of eight for GWs sourced from both vector (solid) and scalar (dashed) boson clouds. The contour lines correspond to the strain sensitivities of aLIGO (blue), tuned GEO600 (orange), and GEO600 with different MSR detuning angles. The shaded gray region corresponds to the parameter space where the superradiance condition fails. An SNR of eight can be achieved at frequencies as large as $31$ kHz, with the highest frequency obtained at a detuning angle of $\phi=45\degree$ (not plotted). Right: SNR of GWs sourced from both vector (solid) and scalar (dashed) boson clouds across a range of boson rest energies for a 0.3 $M_{\odot}$ BH. The SNR curves are calculated with the same strain sensitivities as the left plot and have the same corresponding colors. The dotted black line represents where the SNR is eight. The sharp increase in SNR in aLIGO at 37.5 kHz (FSR) arises due to sideband resonance in its Fabry-Perot cavities, while the kink in the SNR function on the right plot around 14 kHz corresponds to the sudden drop-off in integration time when $\tau_{\text{GW}}$ falls below one week. All GWs are sourced from within the galaxy, being simulated from a distance of $30$ kPc, and assume a sky-averaged interferometer response.
@@
where $\omega_{\mu}=\mu/\hbar$, $m=1$, and $\Omega_{\mathrm{BH}}(0.7)=1.38\times 10^{5}\!\left(\!\frac{0.3M_{\odot}}{M}\!\right)\text{s}^{-1}$. Rearranging this superradiance condition constrains the rest energy of bosons surrounding a BH of mass $M$ to be $\mu<9.10\times 10^{-11}\!\left(\!\frac{0.3M_{\odot}}{M}\!\right)\text{eV}$. For boson clouds with boson rest energies in the target range of $10^{-12}$ to $10^{-10}$ eV, the GW emission frequency is given by [86]
$f_{\text{GW}}\approx 50\,\text{kHz}\times\frac{\mu}{10^{-10}\text{ eV}}.$ (6)
@@
$\hskip-7.59995pth_{0}^{v}\approx 10^{-22}\!\left(\!\frac{M}{0.3M_{\odot}}\!\right)\!\!\left(\frac{\alpha}{0.2}\right)^{\!5}\!\!\left(\!\frac{0.03\,\text{Mpc}}{r}\!\right)\!\!\left(\!\frac{0.7-\chi_{f}}{0.1}\!\right)\!,$ (7)